10. The equation for the velocity of an airplane slowing to a stop to the East on an airway is

𝑣⃗(𝑑) = 70π‘š/𝑠²

𝑑 βˆ’ 44π‘š/𝑠³ 𝑑² . Derive the formula for its instantaneous acceleration and find the velocity when t=1s.


1
Expert's answer
2021-10-20T10:40:36-0400

The instantaneous acceleration:


"a=\frac{\text dv}{\text d t}=70-88t.\\\space\\
v(1)=70·1-44·1^2=26\text{ m/s}."